OpenStudy (anonymous):

PLEASE HELP WITH A COUPLE QUESTIONS!!! WILL GIVE WHATEVER THE HECK IT IS Malloy solved the equation −5x − 16 = 8; his work is shown below. Identify the error and where it was made. −5x − 16 = 8 Step 1: −5x − 16 + 16 = 8 + 16 Step 2: −5x = 24 Step 3: negative five times x over five = twenty-four over five Step 4: x = twenty-four over five Step 1: He should have subtracted 16 from each side. Step 1: He should have divided both sides by −5. Step 3: He should have multiplied both sides by −5. Step 3: He should have divided both sides by −5.

OpenStudy (anonymous):

−5x − 16 = 8

OpenStudy (anonymous):

Mallow solved the equaiton -5x - 16 = 8

OpenStudy (anonymous):

\[−5x - 16 + 16 = 8\]

OpenStudy (anonymous):

-5x = 24

OpenStudy (anonymous):

flutter it nvm

OpenStudy (anonymous):

to much stuff u dont need to answer :/

OpenStudy (anonymous):

all my questions are equations

OpenStudy (anonymous):

\[-1/3.(x − 15) = −4\] Multiply each side by negative one over three , subtract 15 from each side Multiply each side by −3, add 15 to each side Multiply each side by negative one over three , add 15 to each side Multiply each side by 3, subtract 15 from each side

OpenStudy (anonymous):

see

OpenStudy (anonymous):

1 \ 8x (8x + 15) = 24\] 9 49 one hundred seventy-seven over eight two hundred seven over eight

OpenStudy (anonymous):

no idea its fine

OpenStudy (anonymous):

he should have divided both sides by -5